ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 148, 2005

Aaronson, J. and H. Nakada On the mixing coefficients of piecewise monotonic maps

1

Abels, H., G. A. Margulis and G. A. Soifer The Auslander conjecture for groups leaving a form of signature (n−2, 2) invariant

11

Bergelson, V. Multiplicatively large sets and ergodic Ramsey theory  23

Cohen, G. and M. Lin Extensions of the Menchoff–Rademacher theorem with applications to ergodic theory

41

Comets, F. and O. Zeitouni Gaussian fluctuations for random walks in random mixing environments

87

Cranston, M. and S. Molchanov Limit laws for sums of products of exponentials of iid random variables

115

Danilenko, A. I. and M. Lemańczyk A class of multipliers for W^{⊥} 137

Einsiedler, M. and A. Katok Rigidity of measures—The high entropy case and non-commuting foliations

169

Ellis, D. B. and R. Ellis An icer approach to the theory of minimal flows

239

Ellis, R. See D. B. Ellis

Farkas, H. M. and Y. Godin Logarithmic derivatives of theta functions

253

Frantzikinakis, N. and B. Kra Polynomial averages converge to the product of integrals

267

Glasner, E. Topological weak mixing and quasi-Bohr systems  277

Glasner, E., B. Tsirelson and B. Weiss The automorphism group of the Gaussian measure cannot act pointwise

305

Godin, Y. See H. M. Farkas

Goldsheid, I. Ya. and B. A. Khoruzhenko The Thouless formula for random non-Hermitian Jacobi matrices

331

Heath-Brown, D. R., C. E. Praeger and A. Shalev Permutation groups, simple groups, and sieve methods

347

Katok, A. See M. Einsiedler

Khoruzhenko, B. A. See I. Ya. Goldsheid

Kra, B. See N. Frantzikinakis

Lemańczyk, M. See A. I. Danilenko Lin, M. See G. Cohen Margulis, G. A. See H. Abels Molchanov, S. See M. Cranston Nakada, H. See J. Aaronson Praeger, C. E. See D. R. Heath-Brown Shalev, A. See D. R. Heath-Brown Soifer, G. A. See H. Abels Tsirelson, B. See E. Glasner Weiss, B. See E. Glasner Zeitouni, O. See F. Comets